### v1

@@ 2.2.2 Quantum Services Computing
Despite their benefits, QC systems of today lack the capabilities of executing quantum algorithms wrapped with large and diverse amounts of data, inputs, and outputs [2] [10]. As shown in Figure 2 a), large volumes of data in quantum algorithms require more QuBits and complex QuGates that result in deep quantum circuiting and consequently increased error rate, generally referred to as noisy intermediate-scale quantum (NISQ) [4]. To address the issues like NISQ, the classic-quantum split pattern slices the overall quantum software or application into classical modules (pre/post-processing) and quantum modules (quantum computation) that result in hybrid applications[27]. One of the prime examples of classic-quantum split patterns is Shor’s algorithm which involves quantum computations for finding the prime factors of an integer with its application in computer security and cryptography. Quantum computing vendors such as Amazon, IBM, and Google have started to offer their QC systems and infrastructures to be utilised by individuals and organisations by means of quantum services computing [14]. For example, Amazon Braket allows quantum computing services to build, execute, and simulate quantum software based on pay-per-shot model. Ongoing research and development are focused on offering algorithms, hardware, simulations, and mathematical problems as services for quantum platforms [15].
@@
Classical vs. quantum microservices: In service computing model microservices help software designers and developers to structure software applications, via the application of microservice architecture style, as loosely coupled and deployable modules of code that enable computation and data storage [28]. Microservices systems enable enterprises to adopt the service computing model by migrating or modernising their monoliths to a single application that consists of multiple small services [26]. Industry-leading service providers such as Netflix and Amazon represent typical examples of how microservices architecture helps businesses deliver their core business functionality (e.g., video streaming, online shopping) with increased scalability that efficiently serves millions of customers across the globe. Quantum microservicing is a term that refers to microservices that are deployed and executed on quantum computing platforms. For example, the solution [29] utilises Qiskit an open-source quantum software development kit to develop and deliver quantum computation as a collection of microservices that can be executed on IBM Quantum platform. From an implementation perspective, which primarily involves writing and executing source code, both classical and quantum microservice are similar because they implement an algorithm or contain a module of source code that can be executed on QCs [14]. However, from operational and deployment point of view, which focuses on managing the execution of service on quantum platforms, quantum software requirements need to be fulfilled as implemented functionality and desired quality of the microservice.
@@
We conclude that, in order to harness QC as utility computing, there is a need to tailor existing principles and methods of service-orientation or develop new architectures, frameworks, and empirically grounded processes to synergise QC and SOA in the context QCaaS. QSE can provide the conceptual foundations, architectural blueprints, and reusable patterns necessary to navigate the complexities of QCaaS. By embracing SE principles—such as the quantum-classical split, architectural mapping to quantum circuits, and model-driven code generation—developers can orchestrate hybrid applications that synergise classical and quantum computation. Quantum service-orientation when viewedăfrom a utility computing perspective has the potential to bridge the quantum divide — a pressing issue highlighted at the World Economic Forum 2023 — between states and entities that possess quantum computing systems, technologies, and infrastructure, and those that do not [30]. The QCaaS also provides foundations for building a quantum computing cloud (QC cloud or simply quantum cloud), led by technology giants like IBM to offer computation resources to the end-users based on cloud computing [31].
@@ 3 Research Method
  • Architecture-based Development (ABD): As part of SE process, ABD is also referred to as architecture-centric software development [36] where architecturally significant requirements [37] are modelled as system blue-print to guide the subsequent phases of ABD, such as the conceptualisation, implementation, and evaluation of software to be engineered. ABD is generally classified into three main activities, namely architectural analysis, architectural synthesis, and architectural evaluation [38]. The first activity, architectural analysis, is driven by a scenario or architecture story as a description of the problem or potential use-cases for the architecture. The next activity architectural synthesis consolidates the details in the story to create an architecture model or representation that can act as a point of reference, visualising the structural (de-) composition of the software. Finally, the last activity architectural evaluation evaluates the synthesised architecture against ASRs based on scenarios from the architectural story.
@@ 3.1 Step 1 - Specifying the Research Questions
Objectives - Emerging trends indicate open challenges or potentially futuristic research that can shape the agenda, provide a road map for progressing research and development, and outline new hypotheses to be tested for QCaaS. We have identified the trends that can help pinpoint the prevalent challenges and their solutions to help researchers and practitioners in developing solutions for emerging and next-generation of QCaaS.
Table 1: Criteria for Screening and Qualitative Assessment of Selected Studies
Study Selection Step I - Screening of Identified Studies
S1 - The study does not discuss any solution or proposal for quantum computing as a service
S2 - The study is not reported in English
S3 - The study is a duplicate study. Duplicate studies are studies with overlapping contents, e.g., a conference paper extended as a journal article.
S4 - The study is a secondary study/survey paper
@@
Study Selection Step II - Quality Assessment of the Identified Studies
Q1 - Study objectives and Contributions are clear? [Yes = 1, Partially = 0.5, No = 0]
Q2 - Research method to conduct the study is reported [Yes = 1, Partially = 0.5, No = 0]
Q3 - Design and/or implementation details of solution are provided [Yes = 1, Partially = 0.5, No = 0]
Q4 - Details for Experiments/Evaluation/Demonstration of Solution are provided [Yes = 1, Partially = 0.5, No = 0]
Q5 - Study limitations and needs for future research are discussed [Yes = 1, Partially = 0.5, No = 0]
Exclude the study that has a quality assessment score (Q1 – Q5) less than 2.0, otherwise
@@
3.2 Step 2 - Search String and Electronic Data Sources
The initial identification of key terms or keywords for literature was carried out by the first three authors through an analysis of the RQs. The goal of identifying the key research terms was to formulate a search string that can be customised as per specific Electronic Data Sources (EDS) for an automated search, as per the guidelines in [40]. Ultimately, the authors reached a consensus on the search string in Listing 1. The key terms were combined using the boolean operators “OR” and “AND” to formulate the search string. This final search string was developed following a pilot search of relevant literature on IEEE Xplore and Google Scholar. The purpose of the pilot search was to uncover existing study titles and various synonyms for software engineering and quantum software engineering in the context of software services. For example, we discovered that using the term ‘framework’ instead of the term ‘process’ returned a significant number of irrelevant studies focused on quantum computing frameworks that were related to QC in general with no relevance to QSE or quantum service computing. Moreover, inclusion of key-terms such as ‘qubit’ retrieved hardware-focused studies, not software and/or service models, thus adding irrelevant studies (noise). We also dropped the key terms ‘quantum–classical’ and ’quantum classic’ as both overlapped with the main term ’quantum’ that alone sufficiently retrieves most relevant studies and avoids an extensive list of irrelevant studies to be scanned.The primary goal of the final search string was to identify the most relevant literature while minimizing the inclusion of irrelevant studies, thereby leading to a manual review of titles, keywords, and abstracts for study selection.
@@ 3.2 Step 2 - Search String and Electronic Data Sources
3.3 Step 3 - Study Selection - Primary Studies for the SMS
To select the primary studies for SMS, we followed the recommendations by Kitchenham et al. [39] for including or excluding identified studies. Table 1 presents a two-step criteria-driven process for literature inclusion and exclusion via screening and quality assessment. The first step, referred to as screening, enabled us to filter out any irrelevant, redundant, or non-English studies. The second step, namely quality assessment, enabled us to do a qualitative assessment of each included study, eliminating those that did not meet the formulated quality criteria. The established criteria effectively filtered the search results produced by the search string.
@@ 3.3.2 Performing Quality Assessment
The selected studies were evaluated using quality assessment criteria from Table 1 to minimise research bias and assess the significance and completeness of each study. These criteria includes five specific assessment questions (Q1 – Q5), against which each primary study was evaluated. A score of 1 was assigned if a study explicitly addressed the specific QA, 0.5 if the QA was partially addressed, and 0 if there was no evidence of adressing the QA. The total quality score for each study was calculated by summing up the scores assigned for all the QAs. The first author conducted the quality assessment, and the second and third authors independently verified the results for accuracy and consistency. Studies with a cumulative QAs score of 1.5 or higher were included in the final list.
3.4 Step 4 - Extracting Data from Primary Studies
A set of data extraction items were defined to extract specific data related to the RQs. These items represent specific data points extracted from each primary study, aligned with the RQs. The template for data extraction is presented in Appendix A - Table 4. The first author conducted a pilot extraction on 10 studies to evaluate the appropriateness of the data items, which were revised based on feedback from the second and third authors. The first three authors collaboratively performed the formal data extraction. The extracted data was synthesised before documenting the results of the SMS. The extracted data is available via [44].
@@ 3.5 Step 5 - Documenting the Mapping Study
The last step involved documenting the results of SMS, representing answers to individual RQs in dedicated sections. Specifically, the results of the SMS documented as demographic details of published research are in Section 4 (answering RQ-1), software engineering solutions are discussed in Section 5 - 6 (answering RQ-2), and emerging trends are highlighted in Section 7 (answering RQ-3).
@@
Supplementary material for the study is provided as part of the extended details of the documentation. Specifically, the extracted data from primary studies - organized according to the data extraction form presented in Appendix A forms the basis for answering the RQs and documenting the SMS, and is made available via [44]. Moreover, a dedicated GitHub repository and a short video demonstration showcasing the architecture-based development of quantum software are accessible at [45] and [46], respectively.
4 Demography of Published Research (RQ-1)
@@
This section overviews the demographic details of published research to answer RQ-1 based on illustrations in Figure 4. Specifically, RQ-1 presents demographic details about the years and types of publications (Section 4.1), type of research (Section 4.2), and application of research solutions to specific domains (Section 4.3) that complement the answers to other RQs. Demographic details can reveal trends such as most recent research, e.g., journal articles [S3][S33] [^1] The notation [SN], where S = study and N = an integer between 1 - 41 such as S1 (first study in Appendix A - Table LABEL:tab:study-list) indicates a unique reference in the list of primary studies selected for the SMS. This notation also helps to distinguish between the selected primary studies and references in the bibliography section of this paper. published in 2024 propose an innovative solution or evaluate existing research to engineer quantum software for enabling qunatum cloud computing and computational optimisation, further elaborated below. Throughout this paper, while answering the RQs, a summary of the core findings and guidelines representing key takeaway message(s) is provided which is denoted with symbol L.
@@
4.1 Frequency and Types of Publications
Analysing the frequency and types can highlight the relative growth and trends of publications that indicate a) number of peer-reviewed published studies (i.e., frequency: publications per year) and b) prominent fora of publications (i.e., publication types: journal article, conference proceedings etc.) on topic under investigation. The illustrations in Figure 4 a) and Figure 4 b) visualise the core findings about frequency and types of publications discussed below. Specifically, Figure 4 a) shows four years including 2021, 2022, 2023, 2024, presented along the horizontal-axis, and number of publications in a given year, represented on vertical-axis. It is vital to mention that Figure 4 indicates no published study prior to the year 2021 and the reason for this can be two-fold including (i) our literature search (Section 3, Figure 3) did not identify any relevant studies before 2021, or (ii) any identified study before 2021 did not pass the quality evaluation criteria (Table 1) for inclusion in the list, i.e., Appendix A. The studies [S5] [S2] published as workshop paper and conference proceedings in 2021 represent one of the earliest published research on experimenting with microservices that can be executed on quantum computing platforms. The most recent research, for example as reported in [S10][S18] published in 2024 as journal articles focuses on architecting software services for QC platforms. The bar graph reveals a significant growth of published research in the year 2023 and 2024 (i.e., 29/41 studies, reflecting 71% approx.) of the selected studies highlighting progressing of research from experimental solutions [S5, S9] to identifying patterns [S16] and creating reference architecture [S10] that support service engineering for QC software.
Figure 4 b) indicates five types of publications, also referred to as publication fora classified as Journal articles (11/41 studies, i.e., 27%), Conference proceedings (19, 46%), Symposium papers (4, 10%), Workshop Papers (5, 12%), and Book Chapters (2, 5%). Conference proceedings represent the most prominent fora of publications with research focused on publications on quantum function as a service [S3], exploring quantum service-oriantation and cloud computing [S20], quantum microservicing for smart systems [S25], and quantum-classic hybrid computing [S32]. A 2021 report by Scopus also highlights emerging research trends in quantum computing, based on Scopus-indexed documents, with a focus on the demographic details of published research in general area of quantum computing systems, software, and services. The report offers a comprehensive overview, analysing the publication focus, leading institutions, and key contributors in the field of quantum computing [47].
@@ 4.2 Types of Research
The type of research, also commonly referred to as research type facet classifies the published studies based on their specific contribution, such as an innovative model, architecture, framework, or prototype as a newly proposed solution [S1, S3, S10] or validation of an existing solution with a new dataset or case studies [S15]. The classification for the type of research is based on six research type facets namely experience report, evaluation research, opinion papers, philosophical papers, solution proposal, and validation based on classification criteria in [21]. Our analysis did not find any experience report that documents personal experiences or perspectives of researchers and practitioners on a specific problem and/or project, shown in Figure 4 c). Solution proposal is the most prominent type of published research (26 studies, 63%) that document solutions such as processes [S26, S33], patterns [S16, S18], architectures [S10], frameworks [S3, S1], and prototype [S29] etc. for quantum software services. In contrast, we only identified one philosophical paper [S8] that argues about hybrid nature of quantum services that split computation as classical and quantum components. The paper also sketches a toolchain to synergise classical and quantum computations into a hybrid computing model. Evaluation research (4 studies), opinion papers (4 studies), and validation research (6 studies) are reported in a total of 14 studies, representing a total of 34% approx. Specifically, evaluation research aims to evaluate an existing solution based on experimentation, such as a case study on quantum service-orientation in smart cities [S11] and evaluating system performance via simulation of distributed quantum computing [S13]. Opinion papers discuss authors’ negative or positive opinions, i.e., pros and cons regarding a specific technique such as orchestration [S18], deployment [S15], and cloud-based [S19] quantum servicing. Finally, the validation research aims to validate the stated hypothesis, objectives, or proposed features of the solution, which have not yet been deployed in a real-world environment. These types of studies focused on validating the integration of quantum-classic microservices [S5], support of quantum algorithm as a service [S6] and enabling the quantum SOA [S7].
L Type of research Analysing the types of published research highlights that solution proposals represent the most prominent type of published research with a total 26 out 41. QCaaS as a recent and continuously evolving phenomena, i.e., subdomain of QSE relies on theory and practices of service-orientation to propose solution for quantum service-orientation. The role of validation and evaluation type research (total of 8 studies) can be vital to empirically assess existing solutions and methods that can be applied to emerging challenges of QCaaS.

### v2

@@ 2.2.2 Quantum Services Computing
Despite their capabilities and advantages, current QC systems struggle to execute quantum algorithms that involve large and diverse datasets [2] [10]. As shown in Figure Figure 2 a), handling large volumes of data requires more QuBits and complex QuGates, resulting in deeper quantum circuits and higher error rates, a challenge commonly referred to as noisy intermediate-scale quantum (NISQ) [4]. To mitigate issues like NISQ, the quantum-classic split pattern divides quantum applications into classical modules (for pre- and post-processing) and quantum modules (for quantum computation), creating qunatum-classic hybrid applications [29]. A notable example of such hybrid computing is Shor’s algorithm [30], which uses quantum computation to find prime factors of integers, with important applications in computer security and cryptography. QC vendors such as Amazon, IBM, and Google now provide access to their QC systems and infrastructure via quantum services [14]. For instance, Amazon Braket allows users to build, execute, and simulate quantum software using a pay-per-shot model. Current research and development continues to focus on delivering algorithms, hardware, simulations, and mathematical problem-solving capabilities as services on quantum platforms [15].
@@
Classical vs. quantum microservices: In the service computing model, microservices allow software designers and developers to structure applications as loosely coupled, deployable modules that handle computation and data storage [31]. This approach enables enterprises to adopt service-oriented models by migrating or modernising monolithic applications into collections of smaller, independent services [28]. Industry-leading service providers like Amazon, Google, and Netflix illustrate how microservices architecture supports core business functions such as video streaming or online shopping, while providing scalable solutions that serve millions of users worldwide. Quantum microservicing is a term that refers to microservices that are deployed and executed on quantum computing platforms. For instance, the solution in [32] uses Qiskit, an open-source quantum software development kit, to implement and deliver quantum computations as a set of microservices that run on the IBM Quantum platform. From an implementation perspective, focused on writing and executing source code, classical and quantum microservices are similar, as both encapsulate algorithms or code modules that can be executed on their respective platforms [14]. However, from an operational and deployment perspective, which involves managing service execution on quantum platforms, quantum microservices must meet specific software requirements, encompassing both the implemented functionality and the desired quality attributes of the service.
@@
We conclude that, in order to harness QC as utility computing, there is a need to tailor existing principles and methods of service-orientation or develop new architectures, frameworks, and empirically grounded processes to synergise QC and SOA in the context QCaaS. QSE can provide the conceptual foundations, architectural blueprints, and reusable patterns to manage the complexities of QCaaS. By adopting SE principles such as the quantum-classical split, mapping of architectural components to quantum circuits, and model-driven code generation, developers can design and implement hybrid applications that synergise classical and quantum computations. From a utility computing perspective, quantum service-orientation has the potential to bridge the “quantum divide” – a challenge highlighted at the World Economic Forum 2023 – between entities that possess quantum computing systems and infrastructure and those that do not [33]. QCaaS also lays the groundwork for quantum computing clouds (QC cloud or simply QCloud), where technology leaders like IBM provide quantum resources to users through a cloud-based model [34].
@@ 3 Research Method
  • Architecture-based Development (ABD): As part of SE process, ABD is also referred to as architecture-centric software development [38] where architecturally significant requirements [39] are modelled as system blue-print to guide the subsequent phases of ABD, such as the conceptualisation, implementation, and evaluation of software to be engineered. ABD is generally classified into three main activities, namely architectural analysis, architectural synthesis, and architectural evaluation [22]. The first activity, architectural analysis, is driven by a scenario or architecture story as a description of the problem or potential use-cases for the architecture. The next activity architectural synthesis consolidates the details in the story to create an architecture model or representation that can act as a point of reference, visualising the structural (de-) composition of the software. Finally, the last activity architectural evaluation evaluates the synthesised architecture against architecturally significant requirements based on scenarios from the architectural story.
@@ 3.1 Step 1 - Specifying the Research Questions
Objectives - Emerging trends indicate open challenges or potentially futuristic research that can shape the agenda, provide a roadmap for progressing research and development, and outline new hypotheses to be tested for QCaaS. The answer can help identify the trends that can help pinpoint the prevalent challenges and their solutions to help researchers and practitioners in developing solutions for emerging and next-generation of QCaaS.
Table 1: Summary of Steps for Screening and Quality Assessment of Studies
Step I - Study Screening
S1 - The study does not propose or discuss any solution for QCaaS
S2 - The study is not published in English.
S3 - The study is a duplicate, such as a conference paper later extended into a journal article (the updated and/or extended version is considered).
S4 - The study is a secondary study or survey paper (reviewed as supplementary material Table 3).
@@
Step II - Quality Assessment
Q1 - Are the study objectives and contributions clearly stated? [Yes:1, Partially:0.5, No:0]
Q2 - Is the research method reported? [Yes, Partially, No]
Q3 - Are design and/or implementation details of the solution provided? [Yes, Partially, No]
Q4 - Are details of experiments, evaluation, or demonstration of the solution provided? [Yes, Partially, No]
Q5 - Are study limitations and future research directions discussed? [Yes, Partially, No]
Exclude any study with a quality assessment score (sum of Q1–Q5) below 2.0; otherwise,
@@
3.2 Step 2 - Search String and Data Sources
The initial identification of key terms or keywords for literature was carried out by the first three authors through an analysis of the RQs. The goal of identifying the key research terms was to formulate a search string that can be customised as per specific Electronic Data Sources (EDS) for an automated search, as per the guidelines in [41]. Ultimately, the authors reached a consensus on the search string in Listing 1. The key terms were combined using the boolean operators “OR” and “AND” to formulate the search string. This final search string was developed following a pilot search of relevant literature on IEEE Xplore and Google Scholar. The purpose of the pilot search was to uncover existing study titles and various synonyms for software engineering and quantum software engineering in the context of software services. For example, we discovered that using the key term ‘framework’ instead of the key term ‘process’ returned a significant number of irrelevant studies focused on quantum computing frameworks that were related to QC in general with no relevance to QSE or quantum service computing. Moreover, inclusion of key terms such as ‘qubit’ retrieved hardware-focused studies, not software and/or service models, thus adding irrelevant studies (noise). We also dropped the key terms ‘quantum-classical’ and ‘quantum classic’ as both overlapped with the main term ‘quantum’ that sufficiently retrieves most relevant studies and avoids an extensive list of irrelevant studies to be scanned. The primary goal of the final search string was to identify the most relevant literature while minimising the inclusion of irrelevant studies, thereby leading to a manual review of titles, keywords, and abstracts for study selection.
@@ 3.2 Step 2 - Search String and Data Sources
3.3 Step 3 - Study Selection for the SMS
To select the primary studies for SMS, we followed the recommendations by Kitchenham et al. [40] for including or excluding identified studies. Table 1 presents a two-step criteria-driven process for literature inclusion and exclusion via screening and quality assessment. The first step, referred to as screening, enabled us to filter out any irrelevant, redundant, or non-English studies. The second step, namely quality assessment, enabled us to do a qualitative assessment of each included study, eliminating those that did not meet the formulated quality criteria. The established criteria filtered the search results produced by the search string.
@@ 3.3.2 Performing Quality Assessment
The selected studies were evaluated using quality assessment criteria from Table 1 to minimise research bias and assess the significance and completeness of each study. These criteria includes five specific assessment questions (Q1 – Q5), against which each primary study was evaluated. A score of 1 was assigned if a study explicitly addressed the specific Q, 0.5 if the Q was partially addressed, and 0 if there was no evidence of adressing the Q. The total quality score for each study was calculated by summing up the scores assigned for all the QAs. The first author conducted the quality assessment, and the second and third authors independently verified the results for accuracy and consistency. Studies with a cumulative Q score of 1.5 or higher were included in the final list.
3.4 Step 4 - Extracting Data from Studies
A set of data extraction items were defined to extract specific data related to the RQs. These items represent specific data points extracted from each primary study, aligned with the RQs. The template for data extraction is presented in Appendix A - Table 4. The first author conducted a pilot extraction on 10 studies (25% approx.) to evaluate the appropriateness of the data items, which were revised based on feedback from the second and third authors. The first three authors collaboratively performed the formal data extraction. The extracted data was synthesised before documenting the results of the SMS - available via [45].
@@ 3.5 Step 5 - Documenting the Mapping Study
The results of the SMS documented as demographic details of published research are in Section 4 (answering RQ-1), software engineering solutions are discussed in Sections 5 - 6 (answering RQ-2), and emerging trends are highlighted in Section 7 (answering RQ-3).
@@
Supplementary material for the study is provided as part of the extended details of the documentation. Specifically, the extracted data from primary studies - organised according to the data extraction form presented in Appendix A provides the basis for answering the RQs and documenting the SMS, and is made available via [45]. Moreover, a dedicated GitHub repository and a short video demonstration showcasing the architecture-based development of quantum software are accessible at [46] and [47], respectively.
4 Demography of Research on QCaaS (RQ-1)
@@
This section overviews the demographic details of published research to answer RQ-1 based on illustrations in Figure 4. Specifically, RQ-1 presents demographic details about the years and types of publications (Section 4.1), type of research (Section 4.2), and application of research solutions to specific domains (Section 4.3) that complement the answers to other RQs. Demographic details can reveal trends such as most recent research, e.g., journal articles [S3][S33] [^1] The notation [SN], where S = study and N = an integer between 1 - 41 such as S1 (first study in Appendix A - Table LABEL:tab:study-list) indicates a unique reference in the list of primary studies selected for the SMS. This notation also helps to distinguish between the selected primary studies and references in the bibliography section of this paper. published in 2024 propose an innovative solution or evaluate existing research to engineer quantum software for enabling qunatum cloud computing and computational optimisation, further elaborated below. Throughout this paper, while answering the RQs, a summary of the core findings and guidelines representing key takeaway message(s) is provided and denoted with symbol L.
4.1 Frequency and Types of Publications
Analysing the frequency and types can highlight the relative growth and trends of publications that indicate a) number of peer-reviewed published studies (i.e., frequency: publications per year) and b) prominent fora of publications (i.e., publication types: journal article, conference proceedings etc.) on topic under investigation. The illustrations in Figure 4 a) and Figure 4 b) visualise the core findings about frequency and types of publications discussed below. Specifically, Figure 4 a) shows four years including 2021, 2022, 2023, 2024, presented along the horizontal-axis, and number of publications in a given year, represented on vertical-axis. It is vital to mention that Figure 4 indicates no published study prior to the year 2021 and the reason for this can be two-fold including (i) our literature search (Section 3, Figure 3) did not identify any relevant studies before 2021, or (ii) any identified study before 2021 did not pass the quality evaluation criteria (Table 1) for inclusion in the list, i.e., Appendix A - Table LABEL:tab:study-list. The studies [S5] [S2] published as workshop paper and conference proceedings in 2021 represent one of the earliest published research on experimenting with microservices that can be executed on quantum computing platforms. The most recent research, for example as reported in [S10][S18] published in 2024 as journal articles focuses on architecting software services for QC platforms. The bar graph reveals a significant growth of published research in the year 2023 and 2024 (i.e., 29/41 studies, reflecting 71% approx.) of the selected studies highlighting a progression of the research from experimental solutions [S5, S9] to identifying patterns [S16] and creating reference architecture [S10] that support service engineering for QC software.
@@ 4.1 Frequency and Types of Publications
Figure 4 b) indicates five types of publications, also referred to as publication fora classified as Journal articles (11/41 studies, i.e., 27%), Conference proceedings (19, 46%), Symposium papers (4, 10%), Workshop Papers (5, 12%), and Book Chapters (2, 5%). Conference proceedings represent the most prominent fora of publications with research focused on publications on quantum function as a service [S3], exploring quantum service-oriantation and cloud computing [S20], quantum microservicing for smart systems [S25], and quantum-classic hybrid computing [S32]. A 2021 report by Scopus also highlights emerging research trends in quantum computing, based on Scopus-indexed documents, with a focus on the demographic details of published research in general area of quantum computing systems, software, and services. The report offers a comprehensive overview, analysing the publication focus, leading institutions, and key contributors in the field of QC [48].
@@ 4.2 Types of Research
The type of research, also commonly referred to as research type facet classifies the published studies based on their specific contribution, such as an innovative model, architecture, framework, or prototype as a newly proposed solution [S1, S3, S10] or validation of an existing solution with a new dataset or case studies [S15]. The classification for the type of research is based on six research type facets namely experience report, evaluation research, opinion papers, philosophical papers, solution proposal, and validation research based on classification criteria in [23]. Our analysis did not find any experience report that documents personal experiences or perspectives of researchers and practitioners on a specific problem and/or project, shown in Figure 4 c). Solution proposal is the most prominent type of published research (26 studies, 63%) that document solutions such as processes [S26, S33], patterns [S16, S18], architectures [S10], frameworks [S3, S1], and prototype [S29] etc. for quantum software services. In contrast, we only identified one philosophical paper [S8] that argues about hybrid nature of quantum services that split computation as classical and quantum components. The paper also sketches a toolchain to synergise classical and quantum computations into a hybrid computing model. Evaluation research (4 studies), opinion papers (4 studies), and validation research (6 studies) are reported in a total of 14 studies, representing a total of 34% approx. Specifically, evaluation research aims to evaluate an existing solution based on experimentation, such as a case study on quantum service-orientation in smart cities [S11] and evaluating system performance via simulation of distributed quantum computing [S13]. Opinion papers discuss authors’ negative or positive opinions, i.e., pros and cons regarding a specific technique such as orchestration [S18], deployment [S15], and cloud-based [S19] quantum servicing. Finally, the validation research aims to validate the stated hypothesis, objectives, or proposed features of the solution, which have not yet been deployed in a real-world environment. These types of studies focused on validating the integration of quantum-classic microservices [S5], support of quantum algorithm as a service [S6] and enabling the quantum SOA [S7].
L Type of research Analysing the types of published research highlights that solution proposals represent the most prominent type of published research with a total 26 out 41 studies. QCaaS as a recent and continuously evolving phenomena, i.e., subdomain of QSE relies on theory and practices of service-orientation to propose solution for quantum service-orientation. The role of validation and evaluation type research (total of 8 studies) can be vital to empirically assess existing solutions and methods that can be applied to emerging challenges of QCaaS.